### v1

# From Neck to Head: Bio-Impedance Sensing for Head Pose Estimation

Abstract.
We present NeckSense, a novel wearable system for head pose tracking that leverages multi-channel bio-impedance sensing with soft, dry electrodes embedded in a lightweight, necklace-style form factor. NeckSense captures dynamic changes in tissue impedance around the neck, which are modulated by head rotations and subtle muscle activations. To robustly estimate head pose, we propose a deep learning framework that integrates anatomical priors, including joint constraints and natural head rotation ranges, into the loss function design. We validate NeckSense on 7 participants using the current SOTA pose estimation model as ground truth. Our system achieves a mean per vertex error of 25.9 mm across various head movements with a leave one person out cross validation method, demonstrating that a compact, line-of-sight–free bio-impedance wearable can deliver head tracking performance comparable to state-of-the-art vision-based methods.
Bio-impedance Sensing, Pose Estimation, Textile Electrode
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06submissionid: 8747ccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paper
1. Introduction
Continuous head pose tracking is essential for applications like AR/VR and human-robot interaction (Murphy-Chutorian and Trivedi, 2010; Liu et al., 2022; Zimmermann et al., 2018). However, existing tracking technologies face challenges for real-world use. Vision-based systems are compromised by occlusion and privacy concerns, while common wearables present their own trade-offs. Inertial Measurement Units (IMUs) suffer from drift, and electromyography (EMG), while effective for gesture recognition systems like the Meta Reality Labs wristband (Meta Reality Labs, 2025), relies on active muscle contractions, making it less reliable for tracking sustained or passive postures (Disselhorst-Klug et al., 2009). This highlights a clear need for a modality that can robustly infer pose from the body’s underlying anatomy, independent of line-of-sight or the nature of muscle activation.
To address this gap, we propose a novel approach using bio-impedance sensing. Our method is fundamentally distinct from and more efficient than the existing Electrical Impedance Tomography (EIT) in pose/gesture estimation study (Zhang and Harrison, 2015; Zhang et al., 2016; Kyu et al., 2024). Whereas EIT aims to reconstruct a full internal conductivity image using complex hardware (often 8+ electrodes) and computationally intensive algorithms, and the pose/gesture is extracted from large number measurements of time-domain information (current and voltage) due to the complex EIT measurement protocol (Dimas et al., 2024), for example, there are 40 channels in EITPose (Kyu et al., 2024), we argue that this is unnecessary for the head pose estimation. Instead, our system bypasses imaging entirely. We use a simpler five-electrode configuration (one common electrode generates stimuli and four measurement channels) to directly measure frequency-domain impedance features (magnitude and phase). The measured impedance is influenced by several factors, including the composition of underlying tissues (e.g., muscle, fat, and skin) (Sanchez and Rutkove, 2017), the geometry and spacing of the electrodes(Briko et al., 2021), and the orientation of muscle fibers or structural deformation (Stupin et al., 2021). This non-tomographic approach allows us to robustly track both dynamic and static poses with significantly reduced hardware and algorithmic complexity.
We proposed the NeckSense: a lightweight necklace integrating soft, dry, and reusable electrodes for comfortable, continuous wear. The hardware is paired with a transformer-based deep learning framework, Imp2Head, which maps the raw impedance signals to 3D head pose. Critically, our model incorporates anatomical constraints, such as natural joint limits, to ensure its predictions are physically plausible and robust. To validate NeckSense’s performance in realistic conditions, we conducted a user study (N=7). The protocol was specifically designed to capture a diverse and representative range of motions found in naturalistic scenarios like conversation and AR/VR interaction, rather than limiting participants to scripted gestures. We used a state-of-the-art vision-based model for ground-truth labeling. Our system achieved a mean per-vertex error of 25.9 mm, a result comparable to the vision-based system’s own inherent error, demonstrating that the Necksense is a competitive alternative.
Overall, our contributions are:
  • We introduce NeckSense, the first wearable system for continuous head pose estimation using multi-channel bio-impedance sensing. It integrates five soft, reusable electrodes into a lightweight, comfortable, and camera-free necklace form factor.
  • We develop a deep learning framework that maps impedance signals to 3D head pose, incorporating biomechanical constraints (e.g., joint limits and rotation ranges) to improve anatomical plausibility and robustness.
  • We validate NeckSense on 7 participants using a camera-based motion capture system as ground truth, achieving a mean per-vertex error of 25.9 mm, comparable to state-of-the-art vision-based methods without requiring line-of-sight.
2. Related Work
Wearable pose estimation typically employs compact sensor modalities, including IMU (Kim et al., 2024; Georgi et al., 2015; Tavakoli et al., 2018; Mollyn et al., 2023; Xu et al., 2024), capacitive sensors (Ziraknejad et al., 2014; Bian et al., 2024; Zhou et al., 2023), bio-transducers, like EMG (Georgi et al., 2015; Aguiar and Bó, 2017; Tavakoli et al., 2018; Meta Reality Labs, 2025; Chen et al., 2021) and acoustical sensors (Iravantchi et al., 2019), due to their small form factor and suitability for body-mounted applications. IMUs estimate head orientation by integrating angular velocity and acceleration signals (Leonardis et al., 2022). Single-IMU setups, typically integrated into headbands or eyeglasses, are simple but susceptible to drift (Severin, 2020). Multi-IMU arrays improve accuracy but add complexity, bulk, and discomfort (Breen et al., 2009). Hybrid IMU-camera approaches partially resolve drift but are heavily influenced by environmental factors such as lighting and visibility. EMG-based systems effectively track gestures and limb movements by capturing muscle activation signals (Georgi et al., 2015; Aguiar and Bó, 2017). However, EMG inherently relies on active muscle contraction, limiting its efficacy for continuous or static head pose estimation. Recent EMG-based head tracking systems necessitate electrodes on neck and facial areas, significantly reducing comfort and social acceptability (Kim et al., 2024). Additionally, variations in electrode placement can severely affect long-term reliability. Capacitive sensing offers another promising alternative for gesture and pose estimation (Bian et al., 2024, 2022), yet its susceptibility to environmental interference limits its robustness in everyday use cases (Bian et al., 2024). Thus, existing wearable sensor solutions face inherent trade-offs among accuracy, portability, usability, and environmental robustness, motivating continued exploration of complementary sensing modalities.
Bio-impedance sensing has been extensively explored for various physiological monitoring tasks, such as respiration (Jeong et al., 2019), hydration (Sato et al., 2017), and swallowing detection (Lin et al., 2019). More recently, bio-impedance has emerged as a promising modality for gesture recognition in human-computer interaction (Waghmare et al., 2023; Zhang and Harrison, 2015; Liu et al., 2025) and activity recognition (Liu et al., 2024b, c, a). Unlike EMG, bio-impedance does not require active muscle contraction, enabling continuous detection of both static and dynamic postures based on tissue geometry changes alone (Sanchez and Rutkove, 2017; Stupin et al., 2021). Furthermore, bio-impedance inherently supports multi-channel volumetric sensing, capturing spatial deformation patterns across body regions to provide richer contextual information than localized EMG or angular IMUs (Zhang and Harrison, 2015).
EIT, a specific bio-impedance approach, typically reconstructs internal conductivity distributions using boundary voltage and current measurements, requiring numerous electrodes (often more than eight) to achieve high-resolution imaging (Zhang and Harrison, 2015; Zhang et al., 2016; Kyu et al., 2024). Traditional EIT employs a pairwise adjacent current stimulation pattern (Sheffield method), resulting in increased hardware complexity and data processing burdens due to the large number of measurements (specifically, $N(N-3)/2$ for an $N$-electrode system). Additionally, EIT’s inverse reconstruction is highly sensitive to noise, often leading to significant artifacts.
In contrast to traditional EIT, our approach, NeckSense, argues against the necessity of full tomographic reconstruction for effective head pose estimation. Our method employs a significantly simpler hardware setup, using one electrode for stimulation and four additional electrodes placed strategically on the neck to measure the response. Instead of reconstructing tomographic images, our system directly leverages frequency-domain features, signal magnitude and phase, to estimate head pose efficiently and robustly. This approach simplifies the system architecture, reduces computational complexity, and enhances resistance to measurement noise, thus distinguishing itself clearly from conventional EIT methods.
By situating NeckSense within existing bio-transducer literature and explicitly clarifying its unique differences from traditional EIT, we highlight the physiological distinctiveness, practical advantages, and novelty of our simplified bio-impedance sensing approach for wearable head pose estimation and human-computer interaction.
3. APPROACH
3.1. General Principle
Bio-impedance sensing is an electrophysiological technique that monitors the conductive properties of biological tissue between pairs of electrodes. Fig. 1 presents the general principle of bio-impedance sensing-based head pose estimation. The measured impedance is influenced by several factors, including the composition of underlying tissues (e.g., muscle, fat, and skin) (Sanchez and Rutkove, 2017), the geometry and spacing of the electrodes(Briko et al., 2021), and the orientation of muscle fibers or structural deformation (Stupin et al., 2021). When the head moves, particularly during yaw (turning left/right), pitch (nodding up/down), or roll (tilting sideways), several physiological changes occur. Muscular contractions and shifts in tissue can alter local conductivity, while deformation of skin and subcutaneous tissue can change the geometry and thickness of the conductive medium. Additionally, slight displacements of the skin may lead to variations in the relative positioning of the electrodes, further modulating the measured impedance. In this work, we attach multiple electrodes around the neck and track changes in bio-impedance to estimate head pose. By leveraging the sensitivity of bio-impedance to anatomical and postural variations, our system enables camera-free, continuous head movement monitoring.
Figure 1. General principle of bio-impedance sensing-based head pose estimation (Five electrodes integrated in the necklace are attached to the neck to monitor the neck impedance variation caused by head movement, the bio-impedance depends on tissue composition, geometry and distance between the electrodes as well as the orientation of muscle fibers or structure deformation)
3.2. Hardware Implementation
To efficiently acquire bio-impedance signals from the neck, we developed a compact, wearable system named NeckSense, as illustrated in Fig. 2. The system comprises two main components: (1) a set of soft, dry electrodes integrated into a neckband, and (2) an impedance data acquisition module.
[…]
The impedance acquisition module consists of two integrated circuits: an Analog front-end (AFE) chip and Microcontroller unit (MCU). AD5941 generates the sinusoidal excitation signal, measures the resulting current, and includes an on-chip Fast Fourier Transform (FFT) accelerator to efficiently compute impedance in the frequency domain. ESP32-S2 interfaces with the AD5941 via SPI, reads the real and imaginary impedance components, converts them to magnitude and phase, and transmits the data via Bluetooth to a host device (e.g., a computer) for further processing. This system design allows continuous, high-resolution tracking of neck bio-impedance in real time, with minimal form factor and power consumption, making it ideal for wearable, daily use.
Since our system does not require image reconstruction, it does not strictly follow the traditional EIT measurement protocol, allowing for flexible electrode configurations. In this study, we selected a five-electrode setup based on considerations of hardware capability, pose estimation performance, and user experience as shown in Fig. 2. The AD5941 chip supports four measurement channels without the need for an external multiplexer. Placing four electrodes at different positions around the neck allows us to effectively capture major head movement directions. While adding more electrodes could potentially enhance performance, it may also degrade user experience and increase system complexity.
Figure 2. NeckSense system Design. The system consists of five soft, dry, and reusable electrodes positioned around the neck, with one electrode serving as a common electrode to generate the stimuli. Impedance is measured across four channels, each capturing both magnitude and phase information. The electrodes require no adhesive, offering improved comfort and reusability. Impedance signals are acquired using the AD5941 analog front-end, controlled by an ESP32-S2 microcontroller. Measurement data is wirelessly transmitted to a computer via Bluetooth for real-time processing.
3.3. Imp2Head: Head Pose Tracking Algorithm
Figure 3. Architecture of the Imp2Head model, a modular encoder-decoder transformer that maps temporal impedance signals to joint rotations. The encoder captures temporal patterns, while the decoder predicts joint angles autoregressively using attention over encoded features. The loss combines mean squared error with biomechanical constraints to ensure anatomically valid predictions.
This work proposes a Transformer-based sequence-to-sequence model to estimate the head pose dynamics of a person using impedance signals. The model predicts the temporal evolution of three anatomical joints: neck, head, and jaw each described using SMPL-compatible axis-angle rotations.
Ground truth joint angles were obtained by processing RGB video using the OS-X SOTA full body pose estimation pipeline (Lin et al., 2023), which extracts full-body SMPL-X parameters for each frame. From this output, we retained only the axis-angle rotations for the neck, head, and jaw joints. All other joints were discarded. To reduce high-frequency jitter and ensure temporal consistency in the SMPL-X joint rotations, we apply a smoothing algorithm to the ground truth trajectories before training. Since axis-angle representations can suffer from discontinuities, we first convert each joint’s rotation from axis-angle to quaternion format. Let $q_{t}\in\mathbb{R}^{4}$ denote the unit quaternion representing the rotation at time $t$. We apply Gaussian smoothing over a temporal window using the following equation:
(1) $\hat{q}_{t}=\frac{\sum\limits_{k=-K}^{K}w_{k}\cdot q_{t+k}}{\left\|\sum\limits_{k=-K}^{K}w_{k}\cdot q_{t+k}\right\|},\quad w_{k}=\exp\left(-\frac{k^{2}}{2\sigma^{2}}\right)$
where $q_{t+k}$ is the quaternion at frame $t+k$, $w_{k}$ are Gaussian weights with bandwidth parameter $\sigma$, $K$ defines the size of the temporal window, The result is normalized to ensure $\hat{q}_{t}$ remains a valid unit quaternion. After smoothing, the quaternions $\hat{q}_{t}$ are converted back to axis-angle format to serve as the final smoothed ground truth joint rotations. This method preserves global rotational structure while effectively removing high-frequency noise from the estimated motion. The model receives as input a sequence of impedance features over a fixed-length window. Specifically, for each time step, impedance measurements are recorded from 4 sensing electrodes, each providing 2 channels (magnitude and phase), yielding an 8 dimensional feature vector, and the input sequence is defined as:
(2) $X\in\mathbb{R}^{B\times L_{\text{in}}\times 8},\quad L_{\text{in}}=9\cdot L_{\text{out}}$
where $B$ is the batch size and $L_{\text{out}}$ is the desired output sequence length. The output is a predicted sequence of SMPL joint rotations:
(3) $\hat{Y}\in\mathbb{R}^{B\times L_{\text{out}}\times 9}$
Each output frame contains 9 continuous values representing the pitch, yaw, and roll angles of the neck, head, and jaw joints.
Fig. 3 shows the architecture of the proposed Imp2Head, which follows a modular encoder-decoder transformer framework, designed to learn temporal mappings between physiological impedance signals and anatomical joint rotations. The architecture begins with a temporal embedding of the input impedance sequence, which transforms raw sensor readings into a learned latent representation. This is achieved by applying a trainable projection layer followed by positional encoding, enabling the model to recognize temporal ordering within the input window. The intuition behind this step is that understanding the sequence and timing of sensor signals is crucial to accurately capture dynamic head and neck movements over time.
The encoded sequence is processed by a stack of self-attention layers in the transformer encoder. These layers capture complex temporal dependencies and contextual interactions across the full input history, allowing the model to form a global understanding of the observed impedance dynamics. This design leverages self-attention to model long-range temporal relationships, reflecting how head and neck movements involve coordinated muscle activations spread over time rather than isolated instantaneous signals.
The decoder operates autoregressively over a learned output query, independent of any ground truth input during inference. It performs cross-attention over the encoder’s latent representation, allowing the decoder to condition future pose predictions on the historical context extracted from the impedance sequence. This enables the model to forecast continuous sequences of joint rotations by conditioning each predicted step on the learned temporal context, mirroring how past movements influence future poses in a time-dependent manner.
Finally, the decoder output is projected back into the joint angle space through a regression head, producing a continuous sequence of 3D joint rotations. This architecture allows for flexible modeling of high-dimensional, time-dependent motion data and supports multi-step forecasting with attention-guided temporal reasoning.
Imp2Head is trained using a loss function that combines mean squared error (MSE) with biomechanical constraints to ensure physically plausible outputs. Let $\theta_{\min}^{(i)}$ and $\theta_{\max}^{(i)}$ represent the anatomical limits for joint $i\in\{1,\dots,9\}$. A penalty term is used to discourage violations of these limits:
(4) $\mathcal{L}_{\text{bio}}=\frac{1}{B\cdot L_{\text{out}}\cdot 9}\sum_{b,t,i}\left[\max(0,\theta_{\min}^{(i)}-\hat{Y}_{b,t,i})^{2}+\max(0,\hat{Y}_{b,t,i}-\theta_{\max}^{(i)})^{2}\right]$
The total loss is given by:
(5) $\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{MSE}}+\lambda\cdot\mathcal{L}_{\text{bio}}$
where $\lambda$ is a tunable hyperparameter controlling the weight of the biomechanical constraint. It ensures the model’s predictions are both temporally coherent and anatomically valid.
4. Evaluation
4.1. Study Design
Table 1. Measured joint rotation ranges (radians) for 7 participants during natural head and neck movements wearing the NeckSense device.(The joint rotation ranges are computed based on the ground truth.)
Person Neck (Pitch, Yaw, Roll) Head (Pitch, Yaw, Roll) Jaw (Pitch, Yaw, Roll)
1 [0.44, 0.50]0.440.50[-0.44,\,0.50][ - 0.44 , 0.50 ], [0.92, 0.89]0.920.89[-0.92,\,0.89][ - 0.92 , 0.89 ], [0.63, 0.68]0.630.68[-0.63,\,0.68][ - 0.63 , 0.68 ] [0.39, 0.47]0.390.47[-0.39,\,0.47][ - 0.39 , 0.47 ], [0.51, 0.49]0.510.49[-0.51,\,0.49][ - 0.51 , 0.49 ], [0.42, 0.48]0.420.48[-0.42,\,0.48][ - 0.42 , 0.48 ] [0.03, 0.48]0.030.48[0.03,\,0.48][ 0.03 , 0.48 ], [0.13, 0.14]0.130.14[-0.13,\,0.14][ - 0.13 , 0.14 ], [0.12, 0.14]0.120.14[-0.12,\,0.14][ - 0.12 , 0.14 ]
2 [0.68, 0.91]0.680.91[-0.68,\,0.91][ - 0.68 , 0.91 ], [0.52, 0.79]0.520.79[-0.52,\,0.79][ - 0.52 , 0.79 ], [0.42, 0.59]0.420.59[-0.42,\,0.59][ - 0.42 , 0.59 ] [0.47, 0.34]0.470.34[-0.47,\,0.34][ - 0.47 , 0.34 ], [0.41, 0.45]0.410.45[-0.41,\,0.45][ - 0.41 , 0.45 ], [0.28, 0.40]0.280.40[-0.28,\,0.40][ - 0.28 , 0.40 ] [0.11, 0.46]0.110.46[0.11,\,0.46][ 0.11 , 0.46 ], [0.09, 0.12]0.090.12[-0.09,\,0.12][ - 0.09 , 0.12 ], [0.07, 0.08]0.070.08[-0.07,\,0.08][ - 0.07 , 0.08 ]
3 [0.97, 0.47]0.970.47[-0.97,\,0.47][ - 0.97 , 0.47 ], [0.77, 0.35]0.770.35[-0.77,\,0.35][ - 0.77 , 0.35 ], [0.62, 0.51]0.620.51[-0.62,\,0.51][ - 0.62 , 0.51 ] [0.25, 0.51]0.250.51[-0.25,\,0.51][ - 0.25 , 0.51 ], [0.47, 0.39]0.470.39[-0.47,\,0.39][ - 0.47 , 0.39 ], [0.40, 0.23]0.400.23[-0.40,\,0.23][ - 0.40 , 0.23 ] [0.06, 0.48]0.060.48[0.06,\,0.48][ 0.06 , 0.48 ], [0.15, 0.04]0.150.04[-0.15,\,0.04][ - 0.15 , 0.04 ], [0.10, 0.09]0.100.09[-0.10,\,0.09][ - 0.10 , 0.09 ]
4 [0.88, 0.93]0.880.93[-0.88,\,0.93][ - 0.88 , 0.93 ], [0.89, 0.98]0.890.98[-0.89,\,0.98][ - 0.89 , 0.98 ], [0.24, 0.51]0.240.51[-0.24,\,0.51][ - 0.24 , 0.51 ] [0.35, 0.28]0.350.28[-0.35,\,0.28][ - 0.35 , 0.28 ], [0.43, 0.45]0.430.45[-0.43,\,0.45][ - 0.43 , 0.45 ], [0.49, 0.46]0.490.46[-0.49,\,0.46][ - 0.49 , 0.46 ] [0.00, 0.50]0.000.50[0.00,\,0.50][ 0.00 , 0.50 ], [0.05, 0.11]0.050.11[-0.05,\,0.11][ - 0.05 , 0.11 ], [0.15, 0.12]0.150.12[-0.15,\,0.12][ - 0.15 , 0.12 ]
5 [0.50, 0.39]0.500.39[-0.50,\,0.39][ - 0.50 , 0.39 ], [1.03, 0.94]1.030.94[-1.03,\,0.94][ - 1.03 , 0.94 ], [0.65, 0.70]0.650.70[-0.65,\,0.70][ - 0.65 , 0.70 ] [0.46, 0.48]0.460.48[-0.46,\,0.48][ - 0.46 , 0.48 ], [0.33, 0.50]0.330.50[-0.33,\,0.50][ - 0.33 , 0.50 ], [0.11, 0.31]0.110.31[-0.11,\,0.31][ - 0.11 , 0.31 ] [0.24, 0.51]0.240.51[0.24,\,0.51][ 0.24 , 0.51 ], $[-0.16,\,-0.01]$, [0.06, 0.12]0.060.12[-0.06,\,0.12][ - 0.06 , 0.12 ]
6 [0.86, 0.82]0.860.82[-0.86,\,0.82][ - 0.86 , 0.82 ], [0.62, 0.22]0.620.22[-0.62,\,0.22][ - 0.62 , 0.22 ], [0.29, 0.41]0.290.41[-0.29,\,0.41][ - 0.29 , 0.41 ] [0.29, 0.43]0.290.43[-0.29,\,0.43][ - 0.29 , 0.43 ], [0.48, 0.27]0.480.27[-0.48,\,0.27][ - 0.48 , 0.27 ], [0.35, 0.37]0.350.37[-0.35,\,0.37][ - 0.35 , 0.37 ] [0.13, 0.46]0.130.46[0.13,\,0.46][ 0.13 , 0.46 ], [0.04, 0.12]0.040.12[-0.04,\,0.12][ - 0.04 , 0.12 ], [0.07, 0.06]0.070.06[-0.07,\,0.06][ - 0.07 , 0.06 ]
7 [1.04, 1.01]1.041.01[-1.04,\,1.01][ - 1.04 , 1.01 ], [0.38, 0.56]0.380.56[-0.38,\,0.56][ - 0.38 , 0.56 ], [0.59, 0.65]0.590.65[-0.59,\,0.65][ - 0.59 , 0.65 ] [0.46, 0.19]0.460.19[-0.46,\,0.19][ - 0.46 , 0.19 ], [0.51, 0.28]0.510.28[-0.51,\,0.28][ - 0.51 , 0.28 ], [0.49, 0.09]0.490.09[-0.49,\,0.09][ - 0.49 , 0.09 ] [0.01, 0.45]0.010.45[0.01,\,0.45][ 0.01 , 0.45 ], [0.06, 0.11]0.060.11[-0.06,\,0.11][ - 0.06 , 0.11 ], [0.15, 0.11]0.150.11[-0.15,\,0.11][ - 0.15 , 0.11 ]
To comprehensively evaluate the performance of the proposed NeckSense system, we conducted a user study involving 7 participants aged 21 to 31 years ($26.45\pm 2.87$, 2 females and 5 males). Each participant was seated comfortably at a table and wore the NeckSense wearable device. Participants were instructed to move their heads freely, performing natural yaw, pitch, and roll movements, and also performing specific movements, including looking down, up, left, right, front-left, and front-right, they were allowed to speak, drink, and use gadgets during the session, to simulate realistic use cases such as conversation or AR/VR interaction. A camera was positioned in front of each participant to continuously record their movements. The resulting video footage was later used to extract the ground-truth head pose for quantitative evaluation. Each participant took part in the experiment for approximately 30 minutes, ensuring sufficient data to analyze a range of head motions and impedance responses under unconstrained conditions. Detailed angle range of each joint per participant can be found on Table 1.
4.2. Experimental Details
Imp2Head is implemented using the PyTorch on a windows system with an NVIDIA RTX 4090 GPU. The model is optimized using the Adam optimizer with an initial learning rate of $1\times 10^{-3}$, and a batch size of 256. A step-based learning rate scheduler is employed, reducing the learning rate by a factor of 0.5 every 25 epochs to aid convergence. We include a biomechanical constraint term in the loss function, weighted by a factor $\lambda=0.1$. The following angle limits (in radians) are imposed for each joint axis (Pitch, Yaw, Roll):
  • Neck: $[-1.05,1.05]$, $[-1.05,1.05]$, $[-0.70,0.70]$
  • Head: $[-0.52,0.52]$, $[-0.79,0.79]$, $[-0.52,0.52]$
  • Jaw: $[0.0,0.52]$, $[-0.17,0.17]$, $[-0.17,0.17]$
To ensure generalization across subjects, we adopt a Leave-One-Person-Out evaluation protocol. This setup ensures that the model is evaluated on completely unseen individuals, which tests its ability to generalize to new users and anatomy. A sliding window approach is used to generate input-output pairs, with 90 frame input windows and 10 frame output sequences.
4.3. Evaluation Metric
To evaluate the accuracy of the predicted pose, two standard metrics are used: Mean Per Joint Position Error (MPJPE) and Mean Per Vertex Error (MPVE), both computed in 3D space after reconstructing the mesh from the predicted pose parameters.
(6) $\text{MPJPE}=\frac{1}{B\cdot T\cdot J}\sum_{b,t,j}\left\|\mathbf{p}_{b,t,j}^{\text{(gt)}}-\mathbf{p}_{b,t,j}^{\text{(pred)}}\right\|_{2}$
[…]
(7) $\text{MPVE}=\frac{1}{B\cdot T\cdot N}\sum_{b,t,n}\left\|\mathbf{v}_{b,t,n}^{\text{(gt)}}-\mathbf{v}_{b,t,n}^{\text{(pred)}}\right\|_{2}$
where $\mathbf{v}_{b,t,n}\in\mathbb{R}^{3}$ is the 3D position of vertex $n$ and $N$ is the number of SMPL mesh vertices (10475).
4.4. Results
Table 2. Comparison of model variants for head pose tracking using SMPL-X. MPJPE and MPVE Metrics are reported in millimeters and computed against pseudo ground truth from OS-X.
Model MPJPE (mm) MPVE (mm)
Neck Head Jaw Avg Neck Head Jaw Avg
Baseline (Ray et al., 2024) 18.2 16.5 15.7 16.8 16.0 14.5 13.9 14.8
Imp2Head 13.0 11.3 10.8 11.7 9.0 8.2 7.9 8.4
Imp2Head + $\mathcal{L}_{\text{bio}}$ 7.5 6.3 6.3 6.7 6.5 5.6 5.6 5.9
Table 2 summarizes the results of the head pose estimation with different models. In this table, the reference ground truth is from the inference of the SOTA vision-based model OS-X, which we denote as the pseudo ground truth. The baseline model built based on convolutional layers(Ray et al., 2024), achieves MPJPE and MPVE scores of 16.8 mm and 14.8 mm, respectively. The proposed transformer-based model Imp2Head significantly reduces these errors to 11.7 mm (MPJPE) and 8.4 mm (MPVE). Further, adding biomechanical constraint terms (Imp2Head + $\mathcal{L}_{\text{bio}}$) leads to the lowest error rates observed, achieving an MPJPE of 6.7 mm and an MPVE of 5.9 mm. These results indicate that our proposed Imp2Head model incorporating anatomical biomechanical constraints significantly improves head pose estimation accuracy.
Table 3. Comparison of related works and Imp2Head (Proposed approach). Corrected error estimates reflect expected deviation from true 3D ground truth.
Model MPJPE (mm) Notes
IMUPoser1 (Mollyn et al., 2023) 105.0 3 IMUs vs MoCap
MobilePoser1 (Xu et al., 2024) 126.0 3 IMUs vs MoCap
PoseMamba (Huang et al., 2025) 37.5 Vision-based PD vs MoCap
OS-X (Lin et al., 2023) 24.9 Vision-based PD vs MoCap
Imp2Head 6.7 Imp2Head vs Pseudo GT (Vision-based PD)
Imp2Head 25.9 Imp2Head vs True GT (MoCap)
  • 1
    The results are from the full body pose estimation. The work (Huang et al., 2025) notes a challenge in accurately tracking the head and neck regions, indicating these areas can have lower performance scores compared to limbs in full-body models
Since the model is trained and evaluated against a vision-based pose estimation model, which itself is not ground truth but an approximation derived from vision-based inference, the reported MPJPE underestimates the true error. To approximate the model’s actual deviation from real 3D ground truth (e.g., MoCap), we apply a Pythagorean error composition, assuming the errors between OS-X and the impedance-based model are independent:
(8) $\text{MPJPE}_{\text{Corrected}}\approx\sqrt{(\text{MPJPE}_{\text{OS-X}})^{2}+(\text{MPJPE}_{\text{Imp2Head}})^{2}}$
In Table 3, we compare our best impedance-based model (Imp2Head + $\mathcal{L}_{\text{bio}}$) against the state-of-the-art vision-based OS-X model (Lin et al., 2023) and other related works (Mollyn et al., 2023; Xu et al., 2024; Huang et al., 2025). Although our model achieves an MPJPE of 6.7 mm against the pseudo-ground truth provided by OS-X, we recognize that OS-X itself has an inherent error of approximately 24.9 mm compared to a true ground truth system like motion capture (MoCap). Using the Pythagorean error composition, which assumes independent errors between the impedance-based method and OS-X, we estimate the corrected MPJPE of our impedance-based model to be approximately 25.9 mm, which is comparable to the state-of-the-art vision-based system OS-X.
These results affirm that bio-impedance sensing provides a competitive, vision-free, and low-complexity alternative to current vision- and IMU-based head pose estimation methods, suitable for practical wearable scenarios.
5. Discussion and Limitation
Figure 4. Head pose estimation from neck impedance signals: The left column displays time-series neck bio-impedance signals, the middle column shows the predicted SMPL-X poses derived from these signals using Imp2Head, and the right column provides reference ground-truth images of the subjects in corresponding poses.
Our evaluation demonstrates the potential of bio-impedance sensing as a robust, non-intrusive method for wearable head pose estimation. Imp2Head significantly reduces the MPJPE and MPVE compared to baseline models, highlighting the advantage of integrating anatomical biomechanics constraints. Unlike traditional vision-based methods, our impedance-based approach is unaffected by occlusions or varying lighting conditions, making it highly suitable for diverse environments and daily use. The corrected MPJPE of 25.9 mm, while slightly higher than the OS-X vision-based model (24.9 mm), showcases the practical viability of bio-impedance as a competitive alternative. In addition, bio-impedance sensing shows promising potential for jaw pose estimation, a challenging task for other modalities (e.g., IMU, EMG, and capacitive sensors), typically limited to neck-mounted placement. As demonstrated in Fig. 4, our neck-impedance-based model accurately tracks both head and jaw poses, closely following the ground truth trajectories. This highlights bio-impedance’s unique capability to capture subtle anatomical movements beyond the scope of conventional wearable sensors.
Despite the promising results, several limitations of this work warrant discussion. First, the system is trained and evaluated using a vision-based pose estimator (OS-X) rather than a gold-standard motion capture (MoCap) system. While convenient, this reliance on pseudo-ground truth introduces potential bias and limits the upper-bound accuracy of our model. Any inaccuracies or drift in the vision pipeline may propagate through training and evaluation, potentially masking failure cases that would emerge with MoCap-based ground truth. Second, although the use of dry, soft electrodes enhances comfort and wearability, the system remains sensitive to minor shifts in electrode placement, caused by neck movement, sweat, or collar adjustment, which can introduce signal noise and drift. This spatial sensitivity may affect robustness in real-world deployments, especially under long-term use or inconsistent device reattachment, underscoring the need for adaptive preprocessing, auto-calibration, or self-alignment mechanisms. Third, the model does not incorporate individual anatomical differences (e.g., neck size, posture habits, or tissue composition), which could affect impedance patterns. Integrating personalized calibration or subject-specific modeling may further enhance performance across diverse users. Furthermore, the limited sample size (N=7) presents a constraint, potentially affecting the robustness and statistical power of the findings.
These limitations suggest several directions for future research. Enhancing personalization and adaptability, optimizing the mechanical and material design of the wearable for long-term comfort, and extending the system to support complex or high-dynamic activities represent key opportunities. Additionally, incorporating true MoCap-based ground truth in future evaluations would provide a more rigorous assessment of model fidelity.
6. Conclusion
In this work, we introduced NeckSense, a novel bio-impedance-based wearable system designed for continuous and unobtrusive head pose estimation. NeckSense leverages multi-channel impedance measurements with soft, dry and reusable electrodes integrated into a lightweight, necklace-style device. Our Transformer-based Imp2Head model effectively captures the relationship between impedance signals and head pose, enhanced by incorporating biomechanical constraints. Experimental evaluation demonstrates that our approach achieves accuracy comparable to state-of-the-art vision-based systems, highlighting the effectiveness and practicality of bio-impedance sensing as an alternative to traditional methods. Future research directions include addressing individual anatomical variability and extending the method to more dynamic motion scenarios. Overall, NeckSense represents a significant advancement in wearable sensing technology, opening new avenues for unobtrusive, reliable, and versatile human pose tracking applications.